$SPAD/src/input bug10069.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool bug10069.output
)set message test on
)set message auto off
)clear all
)set break resume
--S 1  of 8
cot(0.0)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   csc: no reciprocal
--R
--R   Continuing to read the file...
--R
--E 1
--S 2 of 8
csc(0.0)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   csc: no reciprocal
--R
--R   Continuing to read the file...
--R
--E 2
--S 3 of 8
asec(0.0)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   asec: no reciprocal
--R

--R   Continuing to read the file...
--R
--E 3
--S 4 of 8
acsc(0.0)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   acsc: no reciprocal
--R
--R   Continuing to read the file...
--R
--E 4
--S 5 of 8
asech(0.0)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   asech: no reciprocal
--R
--R   Continuing to read the file...
--R
--E 5
--S 6 of 8
acsch(0.0)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   acsch: no reciprocal
--R
--R   Continuing to read the file...
--R
--E 6
--S 7 of 8
coth(0.0)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   csch: no reciprocal
--R
--R   Continuing to read the file...
--R

--E 7
--S 8 of 8
acoth(0.0)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   acoth: no reciprocal
--R
--R   Continuing to read the file...
--R
--E 8
)spool
)lisp (bye)

———-

References^{[1] nothing}

6